# Necessary and Sufficient Conditions for Capacity-Achieving Private Information Retrieval with Adversarial Servers

### v2

@@ 4.4 Wang’s Colluding PIR [6]
$Q_{j}^{(m)}=\begin{pmatrix}1&z_{j}&z_{j}^{2}&\dots&z_{j}^{S-1}\end{pmatrix}\mathcal{M}$ (23)
@@ 5.1 Conditions for Correctness
Theorem 1 (Conditions for Correctness).
@@ 5.2 Conditions for Robust and Byzantine Privacy
Lemma 1 (Sufficient Condition for Privacy).
@@
For any $m,m^{\prime}\in[1:M]$, $q_{j}\in\mathbf{Q}_{j}$, the following equations hold:
@@ 5.3 Conditions for Robust and Byzantine Capacity
Lemma 4 (Conditions for Capacity Achievability).
@@ 5.4 Conditions for Colluding Privacy
For any $m,m^{\prime}\in[1:M]$, $q_{\mathcal{T}}\in\mathbf{Q}_{\mathcal{T}}$, the following equations hold:
@@
$\Leftrightarrow H(\mathcal{Q}_{\bar{\mathcal{T}}}^{(\theta)}|\mathcal{Q}_{\mathcal{T}}^{(\theta)}=q_{\mathcal{T}})=\log M+H(\mathcal{Q}_{\bar{\mathcal{T}}}^{(\theta)}|\theta,\mathcal{Q}_{\mathcal{T}}^{(\theta)}=q_{\mathcal{T}})$ (102)
@@
This result indicates that the number of feasible $F$ must be constant for all $\theta$ even though $T$ queries are obtained by colluding servers, so that it can not infer queries sent to the other servers.
@@ 5.5 Conditions for Colluding Capacity
Theorem 5 (Conditions for Capacity Achievability).
@@ 6.1 Sun’s Colluding and Robust PIR schemes
Sun’s colluding PIR scheme [2] is a capacity-achieving scheme. As shown in (15), the query is constructed so that the decoding matrix $D_{m}$ exists, thereby satisfying the correctness condition. Furthermore, Lemma 1 in [2] shows that MDS encoding does not affect the distribution of the feasible query set of $\frac{T}{S}$ within the entire matrix. Therefore, $T$ colluding servers cannot infer the queries sent to the other servers. This means that the number of feasible $F$ is constant regardless of the desired index $m$, and $|\{F|\mathcal{Q}_{\mathcal{T}}^{(m)}=q_{\mathcal{T}}\}|=|\{F|\mathcal{Q}_{\mathcal{T}}^{(m^{\prime})}=q_{\mathcal{T}}\}|$ holds for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$. Thus, the privacy condition is satisfied. Finally, $\mathbf{e}_{m}$ in (14) is a full-rank matrix. Therefore, $\mathrm{rank}\ Q_{1:S}[\mathcal{I}]=\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}[\mathcal{I}]$ for any $j\in[1:S]$ and for any $\bar{\mathcal{I}}\subset[1:S]\setminus\{m\}$. For any $k\in[1:M]\setminus\{m\}$, $\mathbf{e}_{k}$ are MDS-encoded by $\mathcal{G}(S^{M},TS^{M-1})$. This means that $Q_{\mathcal{T}}[\bar{m}]$ can represent $Q_{1:S}[\bar{m}]$ for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$. Therefore, $\mathrm{rank}\ Q_{1:S}[\bar{m}]=\max_{\mathcal{T}}\ \mathrm{rank}\ Q_{\mathcal{T}}[\bar{m}]$ holds, and the capacity conditions are satisfied.
@@ 6.2 Wang’s Colluding and robust PIR
Wang’s colluding PIR scheme [6] is not a capacity-achieving scheme. As shown in (27), the decoding matrix $D_{m}$ exists, and the correctness condition is satisfied. Any column vector of $\mathcal{M}$ in (23) has $T$ random elements, and only $Q_{j}^{(m)}[m]=\begin{pmatrix}1&z_{j}&z_{j}^{2}&\dots&z_{j}^{S-1}\end{pmatrix}\mathcal{M}[m]$ has a non-random element in addition. However, by collecting only $T$ queries, $Q_{\mathcal{T}}^{(m)}[m]$ has the same realizations as $Q_{\mathcal{T}}^{(m)}[k]$ for any $k\in[1:M]\setminus\{m\}$, because of the $T$ random elements. Therefore, $|\{F|\mathcal{Q}_{\mathcal{T}}^{(m)}=q_{\mathcal{T}}\}|=|\{F|\mathcal{Q}_{\mathcal{T}}^{(m^{\prime})}=q_{\mathcal{T}}\}|$ holds for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$, and the privacy condition is satisfied. Since $Q_{1:S}^{(m)}[m]$ is an $S\times(S-T)$ matrix, $\mathrm{rank}\ Q_{1:S}^{(m)}[m]=S-T$. On the other hand, at least S1S-1 rows of $Q_{1:S}^{(m)}[m]$ are non-zero vectors. Only when $z_{i}=0$ for one of $i\in[1:S]$, and the first row vector of $\mathcal{M}[m]$ is a zero vector, the $i$-th row vector of $Q_{1:S}^{(m)}[m]$ becomes a zero vector. Therefore, $\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}^{(m)}[m]\geq S-1$. Thus, $\mathrm{rank}\ Q_{1:S}^{(m)}[\mathcal{I}]\neq\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}^{(m)}[\mathcal{I}]$ for any $\bar{\mathcal{I}}\subset[1:M]\setminus\{m\}$, and the capacity condition is not satisfied.
@@
This rate is lower than the capacity. In the case of robust PIR, the same discussion applies for any $S$ out of $S+U$ queries; the capacity condition is not satisfied, while the correctness and privacy conditions are satisfied.

### v3

@@ 4.4 Wang’s Colluding PIR [6]
$Q_{j}^{(m)}=\begin{pmatrix}1&z_{j}&z_{j}^{2}&\dots&z_{j}^{S-1}\end{pmatrix}\mathcal{M}$. (23)
@@ 5.1 Conditions for Correctness
Theorem 1 (Condition for Correctness).
@@ 5.2 Conditions for Robust and Byzantine Privacy
Lemma 1 (Sufficient Condition for Robust and Byzantine Privacy).
@@
Let $f\in\mathcal{F}$ be a realization of the random factor $F$. For any $m,m^{\prime}\in[1:M]$, $q_{j}\in\mathbf{Q}_{j}$, the following equations hold:
@@ 5.3 Conditions for Robust and Byzantine Capacity
Lemma 4 (Conditions for Robust and Byzantine Capacity Achievability).
@@ 5.4 Conditions for Colluding Privacy
Let $f\in\mathcal{F}$ be a realization of the random factor $F$. For any $m,m^{\prime}\in[1:M]$, $q_{\mathcal{T}}\in\mathbf{Q}_{\mathcal{T}}$, the following equations hold:
@@
Due to Lemma 5 and Theorem 4, (94) is the sufficient condition for privacy property (5). This result indicates that the number of feasible $f$ must be constant for all $\theta$ even though $T$ queries are obtained by colluding servers, so that it can not infer queries sent to the other servers.
@@ 5.5 Conditions for Colluding Capacity
Theorem 5 (Conditions for Colluding Capacity Achievability).
@@ 6.1 Sun’s Colluding and Robust PIR schemes
Sun’s colluding PIR scheme [2] is a capacity-achieving scheme. As shown in (15), the query is constructed so that the decoding matrix $D_{m}$ exists, thereby satisfying the correctness condition. In this method, $F$ is induced by the randomness of $\mathbf{e}_{1},\dots,\mathbf{e}_{M}$. Lemma 1 in [2] shows that MDS encoding does not affect the distribution of the feasible query set of $\frac{T}{S}$ within the entire matrix. Therefore, $T$ colluding servers cannot infer the queries sent to the other servers. This means that the number of feasible $f$ is constant regardless of the desired index $m$. Thus, $|\{f|\mathcal{Q}_{\mathcal{T}}^{(m)}=q_{\mathcal{T}}\}|=|\{f|\mathcal{Q}_{\mathcal{T}}^{(m^{\prime})}=q_{\mathcal{T}}\}|$ holds for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$, and the privacy condition is satisfied. Furthermore, $\mathbf{e}_{m}$ in (14) is a full-rank matrix. Therefore, $\mathrm{rank}\ Q_{1:S}[\mathcal{I}]=\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}[\mathcal{I}]$ for any $j\in[1:S]$ and for any $\bar{\mathcal{I}}\subset[1:S]\setminus\{m\}$. For any $k\in[1:M]\setminus\{m\}$, $\mathbf{e}_{k}$ are MDS-encoded by $\mathcal{G}(S^{M},TS^{M-1})$. This means that $Q_{\mathcal{T}}[\bar{m}]$ can represent $Q_{1:S}[\bar{m}]$ for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$. Therefore, $\mathrm{rank}\ Q_{1:S}[\bar{m}]=\max_{\mathcal{T}}\ \mathrm{rank}\ Q_{\mathcal{T}}[\bar{m}]$ holds, and the capacity conditions are satisfied.
@@ 6.2 Wang’s Colluding and robust PIR
Wang’s colluding PIR scheme [6] is not a capacity-achieving scheme. As shown in (27), the decoding matrix $D_{m}$ exists, and the correctness condition is satisfied. In this method, $F$ is induced by the randomness of $\mathbf{r}[1],\dots,\mathbf{r}[M]$. Any column vector of $\mathcal{M}$ in (23) has $T$ random elements, and only $Q_{j}^{(m)}[m]=\begin{pmatrix}1&z_{j}&z_{j}^{2}&\dots&z_{j}^{S-1}\end{pmatrix}\mathcal{M}[m]$ has a non-random element in addition. However, by collecting only $T$ queries, $Q_{\mathcal{T}}^{(m)}[m]$ has the same realizations as $Q_{\mathcal{T}}^{(m)}[k]$ for any $k\in[1:M]\setminus\{m\}$, because of the $T$ random elements. Therefore, $|\{f|\mathcal{Q}_{\mathcal{T}}^{(m)}=Q_{\mathcal{T}}^{(m)}\}|=|\{f|\mathcal{Q}_{\mathcal{T}}^{(m^{\prime})}=Q_{\mathcal{T}}^{(m)}\}|$ holds for any $\mathcal{T}\subset[1:S]$ satisfying $|\mathcal{T}|=T$, and the privacy condition is satisfied. Since $Q_{1:S}^{(m)}[m]$ is an $S\times(S-T)$ matrix, $\mathrm{rank}\ Q_{1:S}^{(m)}[m]=S-T$. On the other hand, at least S1S-1 rows of $Q_{1:S}^{(m)}[m]$ are non-zero vectors. Only when $z_{i}=0$ for one of $i\in[1:S]$, and the first row vector of $\mathcal{M}[m]$ is a zero vector, the $i$-th row vector of $Q_{1:S}^{(m)}[m]$ becomes a zero vector. Therefore, $\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}^{(m)}[m]\geq S-1$. Thus, $\mathrm{rank}\ Q_{1:S}^{(m)}[\mathcal{I}]\neq\sum_{j\in[1:S]}\mathrm{rank}\ Q_{j}^{(m)}[\mathcal{I}]$ for any $\bar{\mathcal{I}}\subset[1:M]\setminus\{m\}$, and the conditions for capacity achievability are not satisfied.
@@
This rate is lower than the colluding capacity. In the case of robust PIR, the same discussion applies for any $S$ out of $S+U$ queries; the capacity condition is not satisfied, while the correctness and privacy conditions are satisfied.